### v1

# Constraining the Cosmic-ray Energy Based on Observations of Nearby Galaxy Clusters by LHAASO

Abstract
Galaxy clusters act as reservoirs of high-energy cosmic rays (CRs). As CRs propagate through the intracluster medium, they generate diffuse $\gamma$-rays detectable by arrays such as LHAASO. These $\gamma$-rays result from proton-proton ($pp$) collisions of very high-energy cosmic rays (VHECRs) or inverse Compton (IC) scattering of positron-electron pairs created by $p\gamma$ interactions of ultra-high-energy cosmic rays (UHECRs). We analyzed diffuse $\gamma$-ray emission from the Coma, Perseus, and Virgo clusters using LHAASO data. Diffuse emission was modeled as a disk of radius $R_{500}$ for each cluster while accounting for point sources. No significant diffuse emission was detected, yielding 95% confidence level (C.L.) upper limits on the $\gamma$-ray flux: for WCDA (1–25 TeV) and KM2A ($>$25 TeV), less than (49.4, 13.7, 54.0) and (1.34, 1.14, 0.40) ×$\times 10^{-14}$ ph cm-2 s-1 for Coma, Perseus, and Virgo, respectively. The $\gamma$-ray upper limits can be used to derive model-independent constraints on the integral energy of CRp above 10 TeV (corresponding to the LHAASO observational range $>1\,\mathrm{TeV}$ under the $pp$ scenario) to be less than $(1.96,0.59,0.08)\times 10^{61}\,\mathrm{erg}$. The absence of detectable annuli/ring-like structures, indicative of cluster accretion or merging shocks, imposes further constraints on models in which the UHECRs are accelerated in the merging shocks of galaxy clusters.
LHAASO, Galaxy clusters, $\gamma$-ray source, UHECRs, Coma, Perseus, Virgo
journal: ApJL
1 Introduction
Galaxy clusters, recognized as the largest gravitationally bound entities in the universe, serve as vital cosmological laboratories for investigating a wide range of astronomical phenomena. These clusters typically encompass hundreds to thousands of galaxies, all enveloped in hot, diffuse intracluster gas that extends across several megaparsecs. Radio observations have uncovered non-thermal radio halos and relics within these clusters, indicating the presence of a population of relativistic particles (Feretti et al., 2012).
Cosmic rays within galaxy clusters are generated and accelerated by multiple astrophysical sources. These include supernova-driven galactic winds, active galactic nuclei (AGNs) (Hinton et al., 2007), large-scale intergalactic shocks arising from accretion and merger processes (Colafrancesco & Blasi, 1998; Ryu et al., 2003; Vannoni et al., 2011; Brunetti & Jones, 2014), and turbulent intracluster magnetic fields (Brunetti & Blasi, 2005). The magnetic fields in galaxy clusters, typically on the order of a few $\mu G$, effectively confine cosmic rays, allowing their accumulation (Völk et al., 1996; Berezinsky et al., 1997; Govoni & Feretti, 2004). When these relativistic particles interact with the ambient matter, they can generate detectable high-energy $\gamma$-rays.
Studying galaxy clusters in the $\gamma$-ray band provides crucial insights into the efficiency of particle acceleration, magnetic confinement, and radiation processes involving cosmic rays. Several radiation mechanisms can give rise to $\gamma$-rays in galaxy clusters. The hadronic scenario, which involves inelastic collisions between cosmic ray protons (CRp) and thermal nuclei (Dennison, 1980), is a major potential $\gamma$-ray radiation mechanism in galaxy clusters. Leptonic scenarios can also generate $\gamma$-rays. Relativistic CRp interact with target photon fields, such as the cosmic microwave background (CMB) and infrared background, producing high-energy $\gamma$-rays (Atoyan & Völk, 2000). $\gamma$-ray emission can also arise from IC scattering of secondary electrons, which are produced by Bethe-Heitler processes involving relativistic protons and the CMB photon field, if UHECRs are accelerated within galaxy clusters, as predicted by some models (Kelner & Aharonian, 2008). Another mechanism involves IC radiation emitted by relativistic cosmic ray electrons (CRe), which originate from interactions between very high-energy (VHE) $\gamma$-rays and the diffuse extragalactic background radiation field.
Previous observations of $\gamma$-rays in the MeV to GeV energy range by EGRET (Reimer et al., 2003) provided only upper limits for several galaxy clusters. Similarly, the Whipple telescope reported upper limits at TeV energies for the Perseus and Abell 2029 clusters (Perkins et al., 2006). More recently, stereoscopic instruments such as H.E.S.S. and VERITAS have also set upper limits from TeV observations (HESS Collaboration et al., 2023; HESS Collaboration & Aharonian, 2009; Arlen et al., 2012; Perkins et al., 2006). The analysis of Fermi LAT data by Xi et al. (2018) revealed extended GeV emission from the direction of the Coma cluster, which was later confirmed by Baghmanyan et al. (2022).
Nearby galaxy clusters, owing to their large physical sizes, are anticipated to be extended sources of $\gamma$-rays. This characteristic makes large field-of-view instruments more effective for their study. In this context, LHAASO, with its $\sim$2.3 steradian FoV and unparalleled sensitivity in the VHE-UHE energy range, is highly suitable for such investigations.
In this Letter, we report the LHAASO observations of three nearby galaxy clusters: Coma, Perseus, and Virgo. Leveraging the $\gamma$-ray upper limits obtained from LHAASO observations, we impose stringent constraints on the energy budget of high-energy particles within these clusters above 10 TeV, corresponding to the $\gamma$-ray energy range observed by LHAASO ($>1\,\mathrm{TeV}$) under the $pp$ interaction scenario, in a model-independent manner.
Such constraints can be extrapolated to lower energies to determine the total CRp energy budget of the galaxy cluster and, in turn, the ratio of CRp to thermal energy, $X_{CRp}=\frac{E_{CRp}}{E_{th}}$. Here, $E_{CRp}$ represents the integral energy of CRp above 1 GeV, and $E_{th}$ denotes the total thermal energy. However, it must be noted that this extrapolation is highly dependent on the confinement of CRp within the galaxy cluster and is significantly model-dependent. This aspect will be elaborated on in detail in the subsequent text.
The structure of this paper is as follows. In Sec.2, we present the fundamental background information on the selected clusters. In Sec.3, we conduct an in-depth analysis of the LHAASO-WCDA and LHAASO-KM2A data. Finally, in Sec.4, we discuss the derived cosmic-ray constraints and their astrophysical implications.
2 Facts of the galaxy clusters
Coma, Perseus, and Virgo are the three closest massive galaxy clusters, with redshifts of approximately 0.023 (Giovannini et al., 1993), 0.017 (Hitomi Collaboration et al., 2018), and 0.004 (Mei et al., 2007), and total masses $M_{500}$ of about $6.13\times 10^{14}M_{\odot}$ (Planck Collaboration et al., 2013), $5.77\times 10^{14}M_{\odot}$ (CTA Consortium et al., 2023), and $0.83\times 10^{14}M_{\odot}$ (Simionescu et al., 2017), respectively (as shown in Tab.A.1). Their proximity and large mass render these clusters ideal targets for studying cosmic rays and dark matter.
Observations in other wavelengths have revealed that all three clusters exhibit radio halos, which are a clear indication of a substantial population of relativistic particles. Specifically, the Coma cluster possesses a giant radio halo, which is typically associated with cluster mergers (Cassano et al., 2010). This halo extends from the core of the Coma cluster and features diffuse radio relics near its virial radius (Thierbach et al., 2003). X-ray observations have shown that NGC 4839 is a subgroup in the south-west region of the Coma cluster, connected to its core by a faint X-ray bridge, suggesting a recent merger event (Churazov et al., 2021; Malavasi et al., 2020). The presence of the giant radio halo, in tandem with the merger activities, implies that shock wave dissipation and turbulence resulting from the merging process play significant roles in injecting non-thermal energy into the system (Vannoni et al., 2011).
Perseus and Virgo are well-known as relaxed cool-core (CC) clusters (Peterson & Fabian, 2006), characterized by their dense cores. Perseus showcases a radio mini-halo (RMH) (Gendron-Marsolais et al., 2020, 2021) and X-ray cavities linked to the radio lobes of NGC 1275 (3C84) (Fabian et al., 2000). Similarly, the buoyant bubbles ascending around M87 (Churazov et al., 2001) suggest that active galactic nucleus (AGN) feedback plays a substantial role in Virgo. Moreover, the RMH in M87 is remarkable as one of the earliest discovered radio sources (Bolton & Stanley, 1948). Effelsberg observations uncovered a faint radio halo around M86 in Virgo, indicating a past interaction between M86’s intracluster medium (ICM) and low-density gas at the cluster’s periphery (Vollmer et al., 2004).
In these galaxy clusters, point sources contribute to the $\gamma$-ray emission. During the analysis, it is crucial to distinguish this emission from the extended $\gamma$-ray emission. Fermi observations indicate the presence of three high-energy point sources within the Coma galaxy cluster, which may be associated with radio galaxies or starburst galaxies (Baghmanyan et al., 2022). Other studies have detected $\gamma$-ray emissions from NGC 1275 (Aleksić et al., 2012) and IC 310 (Neronov et al., 2010) in the Perseus cluster, as well as from M87 (Aharonian et al., 2003) in the Virgo cluster. Significantly, several recent publications and alerts from LHAASO have reported detections of M87 (LHAASO Collaboration, 2024a), NGC 1275 (LHAASO Collaboration, 2024b), and IC 310 (Xiang et al., 2024a, b). Thus, we incorporate these sources as point sources in our analysis.
In recent years, multiple detectors have observed and placed constraints on the diffuse $\gamma$-ray emission from these clusters. HESS observations of the Coma cluster have constrained $X_{\text{CRp}}$ to values below 0.2, thereby ruling out the most favorable theoretical models (HESS Collaboration & Aharonian, 2009). Fermi observations have reported the detection of diffuse $\gamma$-ray emission from the Coma cluster, estimating $X_{\text{CRp}}$ to be around 1% (Xi et al., 2018). Subsequent observations have confirmed this finding and provided further detailed analyses (Adam et al., 2021; Baghmanyan et al., 2022). However, these observations face challenges in differentiating between multiple point sources and extended sources, and the spectral indices obtained are relatively soft. Observations at other wavelengths, such as radio observations of the synchrotron emission from secondary pairs in the Coma cluster, yield constraints on $X_{\text{CRp}}$ that range from less than 0.01% to 28%, depending on the assumptions made regarding the magnetic field. For the Perseus cluster, MAGIC telescope observations have constrained $X_{\text{CRp}}$ to less than approximately 1-2% (MAGIC Collaboration et al., 2010, 2016, 2012). Preliminary simulations based on the expected performance of CTA indicate that the constraints on $X_{\text{CRp}}$ could be an order of magnitude stronger than those from previous studies (CTA Consortium et al., 2023). HESS observations of the Virgo cluster have also constrained $X_{\text{CRp}}$ to values below 0.32 (HESS Collaboration et al., 2023). These constraints are predicated on assumptions about the spatial and spectral distribution of cosmic rays and are applicable to different spatial scales.
In this paper, the constraints on $X_{\text{CRp}}$ are uniformly imposed within the $R_{500}$ range for each cluster. The coordinates of the cluster centers for Coma, Perseus, and Virgo are given as (R.A., Dec.): (194.95, 27.98), (49.95, 41.51), and (187.70, 12.39), respectively. The Region of Interest (ROI) is defined as a circular area with a 3-degree radius, centered on the core positions of the three galaxy clusters. The $R_{500}$ values for these clusters, along with other physical parameters, are presented in Table A.1.
In the TeV energy range, $\gamma$-ray flux is attenuated due to pair production ($\gamma\gamma\rightarrow e^{+}e^{-}$) of $\gamma$-rays interacting with background photons from both the CMB and extragalactic background light (EBL). Appendix B provides detailed information on EBL absorption within these clusters.
3 Data analysis
3.1 Analysis method
Table 1: Fitting Results of Coma, Perseus and Virgo cluster by WCDA and KM2A.
Spatial model TS TSdiska Flux U.Lb.
Coma WCDA disk-2.3 6.25 6.25 61.4
disk-2.7 7.33 7.33 69.0
disk-2.3+1pt 13.3 1.51 49.4
disk-2.7+1pt 14.5 2.75 57.1
KM2A disk-3.0 3.51 3.51 1.34
Perseus WCDA disk-2.3+2pt 19.3 0.00 13.7
disk-2.7+2pt 16.5 0.00 13.7
KM2A disk-3.0 1.08 1.08 1.14
Virgo WCDA disk-2.3+1pt 31.72 0.43 54.0
disk-2.7+1pt 31.47 0.19 44.9
KM2A disk-3.0 0.53 0.53 0.40
a The increase in the TS value resulting from the inclusion of the disk model
under the assumption of the existence of other sources in the Spatial model.
b 95% C.L. Flux U.L. of the disk component. (For WCDA, the integration energy range is 1-25 TeV,
while for KM2A, it is above 25 TeV. The unit is 10-14 photon cm-2 s-1.)
LHAASO represents a pioneering, ground-based $\gamma$-ray and cosmic-ray detector array of the new generation. With its broad bandwidth, wide field of view, and all-weather adaptability, it enables comprehensive detection of extended emissions, including those from galaxy clusters.
LHAASO consists of two detectors for $\gamma$-ray detection. The lower-energy data ($E<25\,\text{TeV}$) are collected by the Water Cherenkov Detector Array (WCDA), which covers an area of $78,000\,\text{m}^{2}$. The data collection period spans from March 5, 2021, to July 31, 2023, with a total effective observation time of 784 days. The high-energy $\gamma$-ray data ($E>10\,\text{TeV}$) are recorded by the Kilometer Square Array (KM2A), which consists of 5195 electromagnetic particle detectors (EDs) and 1188 muon detectors (MDs). The KM2A data incorporate observations from the half-array phase, through the three-quarter-array phase, to the full-array phase, covering the period from December 27, 2019, to January 31, 2024, with a total effective observation time of 1389.5 days.
After data selection, the number of events used in this analysis is $4.69\times 10^{9}$ for WCDA and $1.90\times 10^{9}$ for KM2A. For more details regarding the array and the reconstruction of WCDA and KM2A, refer to (Aharonian et al., 2021a, b).
Figure 1: The upper panel shows the WCDA skymaps (1-25 TeV) with the contribution of known point sources subtracted, while the lower panel displays the KM2A skymaps ($>$ 25 TeV). From left to right, they correspond to the Coma, Perseus, and Virgo clusters. The $R_{500}$ radius of these clusters is respectively indicated by gray dashed lines. The known $\gamma$-rays point sources NGC 1275 and IC 310 in the Perseus cluster are respectively marked with cyan squares and upward triangles, while M87 in the Virgo cluster is indicated by a green star.
The WCDA leverages the number of triggered PMT units per event, denoted as $N_{\text{hit}}$, to approximate the event energy. All events are partitioned into six $N_{\text{hit}}$ bins: 60 - 100, 100 - 200, 200 - 300, 300 - 500, 500 - 800, and $>$ 800.
For the KM2A, its datasets are divided into two energy bins per decade, each with a bin width of $\Delta\log_{10}(E)=0.2$, based on the reconstructed energy. Given EBL absorption and uncertainties in EBL models, the use of higher-energy bins becomes unnecessary.
Based on the reconstructed direction of each event, all events are binned into skymaps with a pixel size of $0.1^{\circ}\times 0.1^{\circ}$. The cosmic-ray background is estimated using the direct integration method (Fleysher et al., 2004).
The three-dimensional (3D) algorithm was employed to simultaneously fit both the spectrum and morphology of the source. In this study, we compute the source significance using ${\rm TS}=2\log\left({\frac{\mathcal{L}_{1}}{\mathcal{L}_{0}}}\right)$. Here, $\mathcal{L}_{1}$ represents the maximum likelihood under the alternative hypothesis we intend to test, while $\mathcal{L}_{0}$ denotes the maximum likelihood under the null hypothesis. According to Wilks’ Theorem (Wilks, 1938), the TS value follows a chi-squared distribution with degrees of freedom equal to the difference between those of the null and alternative hypotheses.
To determine the significance in the sky map for this work, we assume that the source is a point source. For WCDA in the energy range of 1-25 TeV, we adopt a power-law spectrum with an index of 2.7; for KM2A at energies above 25 TeV, we use an index of 3.5. In each pixel, the flux is the only free parameter. In line with Wilks’ Theorem, we consider ±$\pm\sqrt{\rm TS}$ as the measure of significance.
To distinguish between the point-source and extended components within these regions, this study employs a likelihood-fitting approach. For the extended radiation, we describe it using a disk centered at the core of each cluster, with a radius equal to its respective $R_{500}$ value. Regarding the energy spectrum of the disk model, we adopt a power-law model with EBL absorption, as presented in the following equation:
$f(E)=A\cdot\left(\frac{E}{E_{0}}\right)^{\Gamma}\cdot e^{-\tau_{ebl}(E,z)}$ (1)
Here, we fix the intrinsic spectral index of the extended radiation. The uncertainty in the spectral index of the extended radiation gives rise to systematic errors. To address this, we test spectral index assumptions of -2.3 and -2.7. Additionally, since LHAASO has reported detections of M87, NGC 1275, and IC 310, we incorporate these sources as point sources in our analysis, using the same spectral model assumption as described in Equation 1. For the hotspot within the Coma cluster, we evaluate models both with and without a point-source component to assess systematic uncertainties, fixing the intrinsic index to -2.3.
3.2 Results
Figure 2: Left panel: The 95% C.L. flux upper limits for the extended emission within the $R_{500}$ radius of the three galaxy clusters, after correction for the EBL, are presented. The darker-blue and lighter-blue arrows represent the data from WCDA and KM2A, respectively. The gray solid and dashed lines represent the $\gamma$-ray fluxes scaled to the strongest constraint for the IC process of secondary electrons produced by the $p-\gamma$ interactions of accelerated cosmic rays with linear and nonlinear DSA scenarios, respectively, as shown in Figs. 7 and 13 of Vannoni et al. (2011). The blue dashed line represents the $\gamma$-ray flux scaled to the strongest constraint for hadronic processes. From top to bottom, the corresponding clusters are the Coma, Perseus, and Virgo clusters. Right panel: The 95% C.L. upper-limit constraints on $X_{\text{CRp}}$ for these three galaxy clusters, under the hypothesis of different CRp spectral indices, are shown. The sequence of galaxy clusters from top to bottom is consistent with the order presented in the left panel. The solid squares in the first plot on the right panel represent the strongest constraints (99% C.L.) previously obtained by HESS (HESS Collaboration & Aharonian, 2009). The hollow pentagrams and crosses represent VERITAS’s and Fermi’s strongest constraints (99% C.L.) on $X_{\text{CRp}}$ within the virial radius of Coma (Arlen et al., 2012). The hollow circles in the second plot on the right denote the 95% C.L. constraints on $X_{\text{CRp}}$ within the virial radius of the Perseus cluster, as determined by MAGIC (MAGIC Collaboration et al., 2016). The hollow squares in the bottom-most plot represent the 99.7% C.L. constraints on $X_{\text{CRp}}$ within the Virgo cluster at $\sim 20\,\text{kpc}$, as determined by HESS (HESS Collaboration et al., 2023). In the bottom two panels, the open crosses represent the 95% C.L. upper limits provided by Fermi (Ackermann et al., 2010).
The fitting results are presented in Table 1. The significance maps of the Coma, Perseus, and Virgo clusters, with all point sources subtracted, for the WCDA and KM2A are shown in Figure 1. Under all model assumptions, we did not detect any significant extended radiation in these clusters. Consequently, we obtained the 95% C.L. upper limits on the extended radiation of these clusters after EBL correction, using likelihood scans, as shown in the left panel of Figure 2. The darker blue and lighter blue represent the upper limits obtained by WCDA and KM2A, respectively.
For the WCDA data, the consideration of point sources has an impact of approximately 2% on the strongest-constraint upper limit of the Coma cluster, which plays a dominant role in the constraints on $X_{\text{CRp}}$. The assumptions of different spectral indices introduce impacts of about 7%, 4%, and 10% on the strongest-constraint upper limits for the Coma, Perseus, and Virgo clusters, respectively.
Due to the negligible impact of the two aforementioned uncertainties on the constraints, subsequent analyses will uniformly adopt a spectral-index assumption of $\Gamma=-2.3$ and will invariably consider point sources to derive the upper limits of the disk component for these galaxy clusters. For the KM2A, since there are no significant point sources or hotspots in the ROI, we only consider a disk component with a fixed spectral index of $\Gamma=-3.0$.
4 Discussion and Conclusion
Based on the aforementioned constraints on $\gamma$-rays, we can place limits on the energy budget of CRp in galaxy clusters. LHAASO measured the $\gamma$-ray emissions above $1\leavevmode\nobreak\ \rm TeV$, which correspond to the CRp energy of more than $10\leavevmode\nobreak\ \rm TeV$ (Kafexhiu et al., 2014; Kelner et al., 2006). Thus, the $\gamma$-ray upper limits we derived above can be used to derive the model-independent CRp energy budget in this energy range directly, which is listed in Table 2.
The CRp are accelerated by various shocks and turbulent environments within galaxy clusters and then diffuse throughout the cluster. The large scale of galaxy clusters and their magnetic fields (on the order of a few $\mu$G) endows CRp with an escape timescale comparable to cosmological timescales. Furthermore, due to the low density of the ICM, the cooling timescale for CRp exceeds the Hubble time, signifying that CRp in galaxy clusters retains information about their acceleration and propagation. The total CRp energy in the cluster reflects the efficiency of cosmic-ray acceleration. Therefore, if we assume that the extended $\gamma$-ray radiation from galaxy clusters is due to the collision of CRp with the ICM, the upper limits previously derived can constrain the physics related to CRp acceleration and propagation for these three galaxy clusters.
In scenarios where CRs accelerated within the cluster experience full confinement, the resulting CRp population would continuously accumulate over time rather than reaching a stationary state. Under the assumption of energy-independent hadronic interaction cross sections, the resulting $\gamma$-ray emission spectrum would then reflect this time-integrated CRp population. The energy budget integrated over the entire energy spectrum of CRp can be used to derive $X_{CR}$. However, the confinement of CRs is determined by the properties of turbulence in these regions, which can result in an energy-dependent confinement time (Blasi et al., 2007). The confinement can then be effective only below some critical energy $E_{c}$. Below $E_{c}$, the confinement time is larger than the Hubble time, and the $\gamma$-ray spectrum is the same as the injected/accelerated one. For energies above $E_{c}$, the $\gamma$-ray spectrum is softened due to the escape of higher-energy CRs. In LHAASO observations, the most stringent constraints come from the first several energy bins of WCDA of several $\rm TeV$, which correspond to the CRp energy of dozens of TeV. Thus, if $E_{c}>30\leavevmode\nobreak\ \rm TeV$, the constraints derived from LHAASO can be extrapolated to lower energies according to the acceleration spectrum, which is typically a power law. On the other hand, if $E_{c}<30\leavevmode\nobreak\ \rm TeV$, we expect a softening of the $\gamma$-ray spectrum in the WCDA energy band. Thus, the CRp energy budget above $10\leavevmode\nobreak\ \rm TeV$ we derived here cannot be extrapolated to lower energies to constrain $X_{CR}$.
In addition to the uncertain confinement conditions, other significant uncertainties persist, particularly concerning the energy and spatial distributions of CRp and the ICM. Under the assumptions that the CRp energy spectrum follows a power-law distribution and that cosmic rays blow 30 TeV can be completely confined within galaxy clusters, when considering the Diffusive Shock Acceleration (DSA) of CRp, we expect the spatially integrated CRp energy spectrum to be the same as the CRp injection spectrum and to be hard. For example, it could have an index of -2.1 or -2.3. However, a low shock Mach number (Gabici & Blasi, 2003) and propagation effects may lead to a softer spectral index in the local spatial region, such as near the CRp source. Thus, we investigated four scenarios with spectral indices of -2.1, -2.3, -2.5, and -2.7. It is important to note that both theoretical analyses and simulation results generally support a hard spatially integrated CRp energy spectrum (index $>$ -2.5) (Vannoni et al., 2011; Pinzke & Pfrommer, 2010). Therefore, the constraints based on an index of -2.1 or -2.3 are likely more realistic and deserve emphasis.
Here, we assume a spherically symmetric distribution for both CRp and the ICM. Hypothesizing that the CRp density distribution adheres to the same form as the ICM density distribution, this is to maintain consistency with similar previous studies (HESS Collaboration & Aharonian, 2009; Arlen et al., 2012; MAGIC Collaboration et al., 2016; HESS Collaboration et al., 2023; Ackermann et al., 2010; CTA Consortium et al., 2023). If we further assume that all CRp are confined within the galaxy clusters, we can derive $X_{CR}$ and list the values in Tab.2. However, it must be noted that the $X_{CR}$ we derived here is highly model-dependent.
This results are also shown in Figure 2. The blue dashed line in the left panel represents the expected $\gamma$-ray flux. This flux is computed using the parametrization of the $\gamma$-ray production cross-section in $pp$ collisions, as derived in (Kafexhiu et al., 2014). The calculation is carried out under the most stringent constraints, assuming a spectral index of $-2.1$. Under this assumption, the Coma, Perseus, and Virgo clusters constrain the value of $X_{\text{CRp}}$ to be less than approximately $0.029$, $0.012$, and $0.04$, respectively. The total thermal energy within the $R_{500}$ radius of the three galaxy clusters, denoted as $E_{th_{500}}$, can be found in Table A.1. Constraints on $X_{CRp}$ under other assumptions of spectral index are presented in Table 2.
Radio observations of the synchrotron radiation of electrons in galaxy clusters have indirectly constrained the energy density of relativistic protons in the Coma galaxy cluster. Assuming a magnetic field of $2\,\mu$G in the galaxy cluster, the constraints on $X_{\text{CRp}}$ are $X_{\text{CRp}}<0.009\%$, $0.01\%$, and $0.07\%$ based on cosmic-ray index assumptions of 2.1, 2.3, and 2.5, respectively (Reimer et al., 2004). Observations of the magnetic field in Coma cluster indicate that the magnetic field at the center of the it ranges from $3.9$ to $5.4\,\mu$G (Bonafede et al., 2010). This implies that the constraints from radio observations are at least an order of magnitude tighter than $\gamma$-ray constraints, highlighting the complementarity of multi-wavelength approaches. However, our $\gamma$-ray constraints remain complementary, particularly for scenarios with suppressed magnetic fields or non-standard secondary pair cooling (Enßlin, 2002; Brunetti et al., 2013). For comparison, the constraints on $X_{cr}$ from other $\gamma$-ray detectors are presented on the right panel of Figure 2.
Table 2: Constraints on the $X_{CR}$ in these three clusters for different assumptions and scenario.
Scenario Assumptionsa $X_{CR}^{b}$ $E_{CRp_{500}}^{c}$ $E_{CRp_{500}}^{d}(>10\text{TeV})$
Coma pp index=-2.1 0.029 $7.77\times 10^{61}$ $1.96\times 10^{61}$
index=-2.3 0.10 $2.89\times 10^{62}$ $1.63\times 10^{61}$
index=-2.5 0.55 $1.49\times 10^{63}$ $1.46\times 10^{61}$
index=-2.7 3.22 $8.66\times 10^{63}$ $1.39\times 10^{61}$
$p\gamma$+IC linear DSA  $\xi=4$ 0.12 $3.23\times 10^{62}$ N/A
nonlinear DSA $\xi=7$ 0.011 $2.96\times 10^{61}$ N/A
Perseus pp index=-2.1 0.012 $2.48\times 10^{61}$ $5.98\times 10^{60}$
index=-2.3 0.045 $9.42\times 10^{61}$ $5.09\times 10^{60}$
index=-2.5 0.24 $4.90\times 10^{62}$ $4.62\times 10^{60}$
index=-2.7 1.16 $2.40\times 10^{63}$ $3.67\times 10^{60}$
$p\gamma$+IC linear DSA  $\xi=4$ 0.082 $1.70\times 10^{62}$ N/A
nonlinear DSA $\xi=7$ 0.008 $1.66\times 10^{61}$ N/A
Virgo pp index=-2.1 0.038 $3.21\times 10^{60}$ $8.33\times 10^{59}$
index=-2.3 0.15 $1.25\times 10^{61}$ $7.30\times 10^{59}$
index=-2.5 0.80 $6.72\times 10^{61}$ $6.81\times 10^{59}$
index=-2.7 4.03 $3.40\times 10^{62}$ $5.59\times 10^{59}$
$p\gamma$+IC linear DSA  $\xi=4$ 0.12 $1.00\times 10^{61}$ N/A
nonlinear DSA $\xi=7$ 0.011 $9.27\times 10^{59}$ N/A
a For the $p\gamma+\text{IC}$ process, here, $\xi$ represents the compression factor
of the shock wave, assuming a shock velocity of 2000 km/s and an upstream magnetic field of 0.3 $\mu G$.
For further details, please refer to the article Vannoni et al. (2011).
b 95% C.L. upper limit for $X_{\text{CR}}$.
c 95% C.L. upper limit for the total energy of CRp within the $R_{500}$ radius. The unit is erg.
d 95% C.L. upper limit for the total energy of CRp with energy $>$ 10 TeV within the $R_{500}$ radius. The unit is erg.
In addition to the $pp$ interaction of CR protons with ambient gas, other mechanisms in galaxy clusters, such as the IC processes of relativistic electrons or the $p\gamma$ processes of UHECR accelerated by accretion/merging shocks, can also generate $\gamma$-ray radiation. Indeed, galaxy clusters are believed to be the most ideal sites for the acceleration of UHECRs, given their large size and thus long-enough confinement time (Hillas, 1984; Vannoni et al., 2011; Zirakashvili & Ptuskin, 2019; Condorelli et al., 2023). These UHECRs rapidly lose energy through Bethe-Heitler pair production interactions with the Cosmic Microwave Background Radiation (CMBR), generating high-energy electron-positron pairs. Subsequently, these electrons quickly lose energy via the IC and synchrotron processes, leaving a $\gamma$-ray signature near the site of UHECR acceleration. In such cases, the emission would follow the spatial distribution of the shocks and present annulus/ring-type morphologies. However, we found no indications of such $\gamma$-ray emissions in the current LHAASO skymaps.
Based on this physical process, in addition to the previous constraints derived from the $pp$ interaction on CRp at tens of TeV, we can set limits on much higher-energy CRs, reaching up to $\sim 10^{18}-10^{19}$ eV. The $\gamma$-ray emission in this scenario reveals a quite hard spectrum, with a peak in the intrinsic spectrum at $10-100\leavevmode\nobreak\ \rm TeV$. This makes LHAASO an ideal instrument for such a study.
In this work, we compared the upper limit we derived with the calculations in Vannoni et al. (2011), where both linear and nonlinear diffusive shock acceleration (DSA) were considered for a fixed total cosmic-ray energy. By comparing Figures 7 and 13 from this study with the upper limits we provided, we further constrained the UHECRs accelerated in galaxy clusters. The solid and dashed gray lines on the left-hand side of Figure 2 respectively show the expected $\gamma$-ray spectra from linear and nonlinear DSA under the aforementioned physical assumptions, scaled to match our most stringent constraints. It should be noted that the upper limit we derived in this work is the intrinsic $\gamma$-ray spectrum before absorption, and it was also compared with the intrinsic $\gamma$-ray spectra calculated in Vannoni et al. (2011).
For linear DSA, we constrained $X_{CRp}$ to be less than 0.12, 0.08, and 0.12 for the Coma, Perseus, and Virgo clusters, respectively. For nonlinear DSA, these limits were tightened to 0.011, 0.008, and 0.011. Detailed information is also provided in Table 2. The tighter constraints in the non-linear DSA scenario are expected, as shock modification produces a harder CR spectrum (Vannoni et al., 2011). As a result, the available kinetic energy transferred to relativistic particles, is accumulated at the highest energies. These high-energy particles dominate the contribution to the pair production and, ultimately, to the IC $\gamma$-ray emissions of the secondary electrons.
It is noteworthy that the propagation effects of UHECRs cannot be entirely disregarded. While Vannoni et al. (2011) did not consider the propagation and escape processes of UHECRs in galaxy clusters, recent simulation studies (Condorelli et al., 2023) show that approximately 40% of $10^{19}\,\text{eV}$ cosmic rays escape from cluster environments with central magnetic fields of $3\,\mu\text{G}$. This significant escape fraction may substantially modify the predicted $\gamma$-ray spectrum. The non-detection of characteristic annulus/ring-type $\gamma$-ray features in LHAASO observations under the $p\gamma$+IC scenario may imply two possibilities: either the UHECR acceleration efficiency is insufficient to reach the required energy densities or UHECRs ($E\gtrsim 10^{18}\,\text{eV}$) have escaped from the cluster environment. We stress that our current conclusions are conditional on the crucial assumption of effective UHECR confinement within galaxy clusters. A comprehensive consideration of particle escape and its energy-dependent effects will be indispensable for future modeling efforts.
To summarize, by leveraging the upper limit of $\gamma$-rays in nearby galaxy clusters, we have derived stringent constraints on relativistic particles above 10 TeV in these systems. Additionally, we also establish stringent upper limit on the contents of UHECRs in these systems, through the $\gamma$-rays generated by secondary electrons via the $p\gamma$ process. The accumulation of LHAASO data in the next few decades will further deepen our understanding of CRs in galactic clusters and the origin of UHECRs.
In this field, multi-messenger observations, such as those involving neutrinos, can also provide important information. Moreover, the absorption of neutrinos is negligible, which gives them a distinct advantage at TeV energies and above. For example, Shi et al. (2023) utilized IceCube data to constrain the average baryon loading factor in galaxy cluster populations. However, due to the current sensitivity limitations of neutrino detectors, these constraints are not yet as stringent as those obtained from $\gamma$-ray observations. The forthcoming next-generation neutrino detectors, such as KM3NET(Margiotta, 2014), ICECUBE Gen-2(Aartsen et al., 2021), Trident(Ye et al., 2023) and HUNT(Huang et al., 2023), will undoubtedly provide more informative observations and offer deeper insights into this field.
We would like to thank all staff members who work at the LHAASO site above 4400 meter above the sea level year round to maintain the detector and keep the water recycling system, electricity power supply and other components of the experiment operating smoothly. We are grateful to Chengdu Management Committee of Tianfu New Area for the constant financial support for research with LHAASO data. We appreciate the computing and data service support provided by the National High Energy Physics Data Center for the data analysis in this paper. This research work is supported by the following grants: The National Natural Science Foundation of China No.12393854, No.12393851, No.12393852, No.12393853, No.12205314, No.12105301, No.12305120, No.12261160362, No.12105294, No.U1931201, No.12375107, No.12173039, the Department of Science and Technology of Sichuan Province, China No.24NSFSC2319, Project for Young Scientists in Basic Research of Chinese Academy of Sciences No.YSBR-061, and in Thailand by the National Science and Technology Development Agency (NSTDA) and the National Research Council of Thailand (NRCT) under the High-Potential Research Team Grant Program (N42A650868).
Appendix A Detailed Information OF THESE GALAXY CLUSTERS
In this appendix, we provide detailed information on the three galaxy clusters, as well as the distribution functions and parameters of $n_{e}$ and electron pressure $P_{e}$ used in this study. All basic information about these three clusters are listed in Table A.1.
To characterize the thermal gas within the ICM across different galaxy clusters, we employed various spatial models for $n_{e}$ and $P_{e}$, as used in previous studies (HESS Collaboration & Aharonian, 2009; Adam et al., 2021; CTA Consortium et al., 2023; HESS Collaboration et al., 2023). Details are provided in Table A.2. For the Virgo cluster, previous studies focused only on the central region ($<$ 44 kpc). In this study, we adopt the latest fitting results for gas and temperature distributions, including data beyond the $R_{500}$ radius, obtained from Planck observations of the Sunyaev–Zel’dovich (SZ) effect (Ade et al., 2016).
For the Perseus and Virgo clusters, the Table A.2 shows the distribution of electron thermal temperature. The pressure distribution can be derived from $n_{e}$ and $T$ using the following equation:
$\centering P_{e}(r)=n_{e}(r)k_{\mathrm{B}}T(r)\@add@centering$ (A1)
Table A.1: The basic properties of Coma, Perseus and Virgo cluster.
Name RA DEC Redshift Size Angular size Mass $E_{th_{500}}^{a}$
(deg) (deg) z $R_{500}$(Mpc) $\theta_{500}$(deg) $M_{500}$($M_{\odot}$) (erg)
Coma 194.95 27.98 0.023 1.35 0.78 $6.13\times 10^{14}$ $2.69\times 10^{63}$
Perseus 49.95 41.51 0.017 1.26 0.96 $5.77\times 10^{14}$ $2.08\times 10^{63}$
Virgo 187.70 12.39 0.004 0.66 2.3 $0.83\times 10^{14}$ $8.42\times 10^{61}$
a Total thermal energy within $R_{500}$ radius.
Table A.2: The spatial distribution models and corresponding parameters for the three galaxy clusters used in this study.
Cluster Physical quantity Function Parameters
Comaa Density profile: $n_{e}(r)=n_{0}\left[1+\left(\frac{r}{r_{c}}\right)^{2}\right]^{-\frac{3\beta}{2}}$ $n_{0}=3.42\times 10^{-3}\mathrm{\leavevmode\nobreak\ cm}^{-3},r_{\mathrm{c}}=290\ \mathrm{kpc},\beta=0.75$
Perssure profile: $n_{e}(r)=\frac{P_{0}}{\left(\frac{r}{r_{p}}\right)^{c}\left(1+\left(\frac{r}{r_{p}}\right)^{a}\right)^{\frac{b-c}{a}}}$ $P_{0}=0.022\ \mathrm{kev}\ \mathrm{cm}^{-3},r_{p}=466.8\ \mathrm{kpc},a=1.8,b=3.1,c=0.0$
Perseusb Density profile: $n_{e}(r)=n_{0,1}\left[1+\left(\frac{r}{r_{c,1}}\right)^{2}\right]^{-\frac{3\beta_{1}}{2}}\\ \quad\quad\quad+n_{0,2}\left[1+\left(\frac{r}{r_{c,2}}\right)^{2}\right]^{-\frac{3\beta_{2}}{2}}$ $n_{0,1}=4.6\times 10^{-2}\ \mathrm{cm}^{-3},r_{c,1}=57\ \mathrm{kpc},\beta_{1}=1.2,n_{0,2}=3.6\times 10^{-3}\ \mathrm{cm}^{-3},r_{c,2}=278\ \mathrm{kpc},\beta_{2}=0.71$
Temperature profile: $k_{\mathrm{B}}T(r)=7\times\left(1+\left(\frac{r}{r_{t,1}}\right)^{3}\right)\\ \times\left(2.3+\left(\frac{r}{r_{t,1}}\right)^{3}\right)^{-1}\left(1+\left(\frac{r}{r_{t,2}}\right)^{1.7}\right)^{-1}\mathrm{keV}$ $r_{t,1}=73.8\ \mathrm{kpc},r_{t,2}=1600\ \mathrm{kpc}$
Virgoc Density profile: $n_{e}(r)=n_{0,1}\left[1+\left(\frac{r}{r_{c,1}}\right)^{2}\right]^{-\frac{3\beta_{1}}{2}}\\ \quad\quad\quad+n_{0,2}\left[1+\left(\frac{r}{r_{c,2}}\right)^{2}\right]^{-\frac{3\beta_{2}}{2}}(r\lesssim 30\ kpc)\\ n_{\mathrm{e}}(r)=\frac{8.5\times 10^{-5}}{(r/\mathrm{Mpc})^{1.2}}\mathrm{\leavevmode\nobreak\ cm}^{-3}(r\gtrsim 250\ kpc)$ $n_{0,1}=8.9\times 10^{-2}\ \mathrm{cm}^{-3},r_{c,1}=4\ \mathrm{kpc},\beta_{1}=1,n_{0,2}=1.9\times 10^{-3}\ \mathrm{cm}^{-3},r_{c,2}=19.4\ \mathrm{kpc},\beta_{2}=0.47$
Temperature profile: $k_{\mathrm{B}}T(r)=\frac{2.4-0.77\times\mathrm{e}^{-(r/\mathrm{Mpc})^{2}/\left(2\sigma_{r}^{2}\right)}}{1+[0.9\times(r/\mathrm{Mpc})]^{2}}\mathrm{keV}$ $\sigma_{r}$ = 23.7 kpc.
  • a
    For Coma, We adopt the $\beta$ model (Cavaliere & Fusco-Femiano, 1978) for the density profile and a generalized Navarro Frenk White (Nagai et al., 2007) model for pressure radial profile, with parameters taken from Briel et al. (1992) and Planck Collaboration: et al. (2013).
  • b
    The functions and parameters used for the Perseus cluster in this study are consistent with those employed in CTA Consortium et al. (2023).
  • c
    Ade et al. (2016) combines a double-beta function for the inner region $(r\lesssim 30\ kpc)$ with a power-law function for the outer region $(r\gtrsim 250\ kpc)$ to describe the thermal electron number density distribution in the Virgo cluster. A simple linear interpolation was applied to the intermediate region. We converted the angular units into distance units.
Appendix B EBL model
In this appendix, we provide detailed information regarding the EBL model. $\gamma$-rays within the energy range observed by LHAASO undergo attenuation due to pair-production interactions with the CMB and EBL as they traverse extragalactic space. The photon-photon pair-production cross-section averaged over directions of the background-radiation field depends on the product of energies of colliding photons. For the given energy of the $\gamma$-ray photon $E_{\gamma}$, it peaks at the wavelength of background photons $\lambda\sim 1(E_{\gamma}/{\rm 1\ TeV})^{-1}\mu\rm m$. Thus, the $\gamma$-ray opacity above and below 100 TeV is due to the absorption on CMB and EBL, respectively. The opacity has been calculated by performing the line-of-sight integral of the product of the pair production cross section with the energy density of the radiation fields. The dependence of the pair production cross section on the energy is given in Gould & Schreder (1967). The EBL energy density is taken from the Franceschini et al. (2008). Figure B.1 illustrates the variation in absorption efficiency with energy at the respective distances of the Coma, Perseus, and Virgo clusters, as predicted by the EBL model adopted in this paper.
Figure B.1: The EBL absorption efficiency for each of the three galaxy clusters. Coma is represented by dotted line, Perseus by dashed line, and Virgo by solid line.